m=7/8, 4,-5)
y-(-5)=7/8(x-4)
y+5=7/8 * 8/8-5
64-5=59
y=7/8x-59

Answer by stanbon(57337)

m=7/8, (4,-5)
y-(-5)=(7/8)(x-4)
y+5=(7/8)x - (7/8)4
---
y+5 = (7/8)x - (7/2)
----
y = (7/8)x - (7/2) - (10/2)
----
y = (7/8)x-(17/2)
==================
Cheers,
Stan H.